Skysong Company has five employees participating in its defined benefit pension plan. Expected years of future service for these employees at the beginning of 2017 are as follows.

Employee

Future Years of Service

Jim 3
Paul 4
Nancy 5
Dave 6
Kathy 6


On January 1, 2017, the company amended its pension plan, increasing its projected benefit obligation by $82,800.

Compute the amount of prior service cost amortization for the years 2017 through 2022 using the years-of-service method, setting up appropriate schedules.

Year

Annual Amortization

2017 $

2018

2019

2020

2021

2022
Answer #1
Jim Paul Nancy Dave Kathy Total
2017 1 1 1 1 1 5
2018 1 1 1 1 1 5
2019 1 1 1 1 1 5
2020 1 1 1 1 4
2021 1 1 1 3
2022 1 1 2
Total 3 4 5 6 6 24
Cost per Service-Year 3450 =82800/24
Year Annual Amortization
2017 17250 =3450*5
2018 17250 =3450*5
2019 17250 =3450*5
2020 13800 =3450*4
2021 10350 =3450*3
2022 6900 =3450*2